OpenStudy (anonymous):

Statistics: The fracture toughness (in ) of a particular steel alloy is known to be normally distributed with a mean of 28.3 and a standard deviation of 0.77. We select one sample of this alloy at random and measure its fracture toughness. Given that the fracture toughness is at least 27, what is the probability that the fracture toughness will be between 29 and 40.3? is it just P(X <= 40.3) - P(X <= 29)? How do I use the 27?

OpenStudy (anonymous):

its on a normal distribution

OpenStudy (anonymous):

This is a problem involving conditional probability. You want to find \[P(29\le X\le40.3~|~X\ge27)\] This means the probability that \(X<27\) is zero, but \(P(29\le X\le40.3)\) alone assumes it's not zero.

OpenStudy (anonymous):

\[\begin{align*}P(29\le X\le40.3~|~X\ge27)&=\frac{P(29\le X\le40.3)\cap P(X\ge27)}{P(X\ge27)}\\\\ &=\frac{P(29\le X\le40.3)}{P(X\ge27)}\\\\ &=\frac{P\left(\dfrac{29-28.3}{0.77}\le Z\le\dfrac{40.3-28.3}{0.77}\right)}{P\left(Z\ge\dfrac{27-28.3}{0.77}\right)}\\\\ &=\frac{P\left(0.91\le Z\le15.58\right)}{P\left(Z\ge-1.69\right)}\\\\ \end{align*}\]

OpenStudy (anonymous):

Thank you! your awesome.